Blank tapes are selling at 3.50 and are reduced at $1. What is the rate of discount?
Answer by palanisamy(496)

Blank tapes rate = $3.50
Reduced rate = $1.
Reduced amount = 3.50-1 = 2.50
Rate of discount = (2.5/3.5)*100 = 71.43%